OpenStudy (anonymous):

Solve –3(2x – 1) = x + 10

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

can you show how you got that please? @MeganLewis

OpenStudy (anonymous):

sure. hold on

OpenStudy (anonymous):

-3(2x-1)=x+10 -6x+3=x+10 -10 -10 -6x-7=x +6x +6x -7=7x divide 7 from both sides so your answer is x=-1

OpenStudy (anonymous):

exactly ^

OpenStudy (anonymous):

thank youuu :)

OpenStudy (anonymous):

You're welcome!